# Scheme-independent determination of the QCD running coupling at all scales from jet observables using the principle of maximum conformality and infinite-order scale setting

### v3

@@ Abstract
We present a new approach to determining the strong coupling $\alpha_{s}(Q)$, over the entire range of validity of perturbative QCD, for scales above $\Lambda_{\mathrm{QCD}}$ and up to the Planck scale $\sim 1.22\cdot 10^{19}$ GeV, with the highest precision and using the data of a single experiment. In particular, we use the results obtained for the thrust ($T$) and $C$-parameter ($C$) distributions in $e^{+}e^{-}$ annihilation at a single annihilation energy $\sqrt{s}=M_{Z}$ (i.e. at the $Z^{0}$ peak). This new method is based on the intrinsic conformality (iCF) and on the Infinite-Order Scale Setting, using the Principle of Maximum Conformality (i.e. the PMC), which allows a rigorous determination of the renormalization scales for the event-shape variable distributions satisfying all of the requirements of Renormalization Group Invariance, including renormalization-scheme independence and consistency with Abelian theory in the $N_{C}\to 0$ limit. This new method is based on the scale-invariance of the iCF, which allows determination of $\alpha_{s}(\mu_{0})$ at any scale $\mu_{0}$, and on the Maximum Likelihood statistical approach. We introduce a novel approach to determining the best-fitting range based on the most-likely-lowest $\chi^{2}$ calculated considering all possible intervals among the entire range of bins available in the perturbative region. This new method improves the precision and leads to results that are statistically more reliable. In particular, using data for thrust and $C$-parameter at the $Z^{0}$ peak from ALEPH, OPAL, DELPHI and L3 experiments, we obtain the average value: $\alpha_{s}(M_{Z})=0.1181^{+0.0005}_{-0.0005}$, for the strong coupling. This determination of $\alpha_{s}(M_{Z})$ is consistent with the world average and has an improved precision with respect to the values obtained from the analysis of event shape observables currently used in the world average.
@@
In this article we shall demonstrate that a recently proposed method [3] for solving the renormalization scale and scheme ambiguities, namely infinite-order scale setting using the principle of maximum conformality (PMC), based on intrinsic conformality (iCF) and by using the PMC, leads to an improved extraction of $\alpha_{s}(Q)$ at the $Z^{0}$ peak from the thrust and $C$-parameter distributions in the annihilation process $e^{+}e^{-}\to 3$ jets. Moreover, our new results agree with the most recent world average given by the Particle Data Group [1], improving the precision and significantly reducing the theoretical errors related to the renormalization scale. In addition, we introduce a new method to enhance the statistics in the Maximum-Likelihood approach, which also reduces the impact of the errors related to hadronization, resummation and other nonperturbative physics.
@@ I The strong-coupling evolution
with the $\beta$-coefficients, $\beta_{n}$, $n=0,...,4$, from one- to five-loop accuracy respectively, calculated in a given scheme. Results for these coefficients can be found in Refs. [39, 40, 41, 42, 43, 44, 45, 46]. For a review of the strong coupling and of all the $\beta$ coefficients, see e.g. [47, 48]. In particular, we use the $\overline{\mathrm{MS}}$ scheme throughout this paper. For the running coupling $\alpha_{s}(Q)$, we use the solutions at 5-loop accuracy given by both the RunDec program [49] and by the perturbative solution given by
@@ II Event-shape variables and the renormalization scale
where $\vec{p}_{i}$ indicates the three-momentum of particle $i$ and the sums run over all particles in the hadronic final state; $\vec{n}$ is a unit vector that is varied to maximize thrust $T$, determining the so-called thrust axis, i.e. $\vec{n}_{T}$. Alternatively, the variable $(1{-}T)$ is often used, which for 3-jet production at LO is restricted to the range $0<(1{-}T)<\frac{1}{3}$. We have back-to-back and spherically symmetric events for $T=1$ and $T=\frac{2}{3}$ respectively. For the $C$-parameter, $\theta_{ij}$ is the angle between $\vec{p_{i}}$ and $\vec{p_{j}}$. For 3-jet production at LO the $C$-parameter is restricted by kinematics to the range $0\leq C\leq\frac{3}{4}$.
Theoretical results for normalized infrared-safe observables, such as the thrust or $C$-parameter distribution for the annihilation process $e^{+}e^{-}\rightarrow 3\,$jets at the $Z^{0}$ peak to NNLO accuracy, are given in Refs. [50, 51, 33, 34, 35, 36, 37]. When using the conventional scale-setting method, only one value of $\alpha_{s}$, at the kinematic scale $\sqrt{s}$ of the process, can be extracted with errors given by the uncertainty in choosing the renormalization scale, while using the PMC, which is based on intrinsic conformality (iCF), it is possible to determine the strong coupling at any scale, even up to the highest scale where the theory is expected to breakdown. As shown in Eq. (I), the final renormalization scale $\mu$ is set by using the PMC method while $\mu_{0}$ can be set to any value in the perturbative range; we can thus determine $\alpha_{s}(\mu_{0})$ at any scale. Results for the PMC for both thrust and $C$-parameter have been discussed in Refs. [2, 52, 3, 53]. We remark that the PMC solves the renormalization scale ambiguity and the scales, related to each order of calculation, are fixed perturbatively by determining the $\beta_{0}$ term of each scale at each order of accuracy. These scales take into account the dynamics of the underlying quark and gluon subprocess and thus fix the magnitude of the strong coupling at each order and for the corresponding value of the selected variable $(C,1{-}T))$. In fact, these scales are not single valued but are themselves functions of the kinematic scale, $\sqrt{s}$, and of the event-shape variable via the coefficients of the number of flavors, $N_{f}$, related to the loop integrals and thus to their UV-divergent part. The PMC scales are fixed order by order and the last scale is set to the kinematic scale of the process $\sqrt{s}\equiv M_{Z}$, given the missing knowledge of the higher-order $\beta_{0}$-term. We have defined this type of uncertainty as the second kind of residual scale dependence [54]. We remark that both the first and second kind of residual scale uncertainties are of the same order as the uncertainties in the first missing higher-order contribution and thus both have a mere perturbative nature linked to the perturbative series and its accuracy.
@@ III Fitting Method
The $\chi^{2}$ analysis has been performed following the approach of Ref. [55]; thus, in the first approximation, considering the bins as being independent of each other and the experimental errors $\sigma_{i}$ determined by summing the statistic and systematic errors (due to detector effects and to hadronization) in quadrature. Thus, we have not considered correlation effects and the errors calculated on the fitted parameter and the same $\chi_{\mathrm{min}}^{2}$-values are considered mainly for sake of comparison among different fitting-regions, different models for the coupling and different methods for the scale-setting; only secondarily these values have been used to determine the goodness-of-fit. In fact, as reported in Ref. [55], without the inclusion of the correlation effects these values can only give a rough measure for the goodness-of-fit.
@@
The high-precision experimental data have been taken from LEP [22]; in particular, these data consist of $42$ experimentally measured values for thrust, $0.420<(1{-}T)<$ 0.42, divided into $42$ bins centered and with spacing $0.01$, while for the $C$-parameter we have $50$ experimental measurements in the range $10<C<$ 1 divided in $50$ bins centered and with spacing of $0.02$. In order to fit the theoretical distributions to the experimental data, for the case of the $C$-parameter we have to reduce the $100$-bin partition shown in Ref. [37] to $50$ bins. This reduction is made by averaging over two consecutive values of the distributions for the LO, NLO and NNLO coefficients.
@@ III.1 Best-fitting region
The purpose of this analysis is to improve or at least to encode the method used in Ref. [4] of setting the fitting interval in the range where: “…  a good perturbative description is available and where hadronization and detector corrections are below …”, in order to enhance the statistics and to eliminate any possible human prejudice or uncertainty at the border of the interval. For the case of the $(1{-}T)$ and $C$-parameter event-shape variables, we know well that in the central region of 3-jet production the perturbative calculations give the best fit to data. Besides, introduction of other corrections, such as nonperturbative corrections and hadronization over the entire range of the event shape, does not improve the theoretical predictions; on the contrary, much larger errors are introduced. Thus, it is a common practice to reduce the fitting region to the perturbative region in order to improve precision and statistics. Although we agree with this procedure, we must admit that it is not entirely consistent with the Maximum-Likelihood procedure adopted to fit the parameters. We are basically adding one degree of freedom – the interval range – and we are eliminating it subsequently by introducing some criteria, which might not be entirely consistent with the minimum $\chi^{2}$ analysis. Here, we now introduce a new procedure to select the BFR interval.
@@
Results of the fits for $\alpha_{s}(M_{Z})$, for each bin interval in the range $0.720.38<C<$ 0.72 (labelled as GFR) are displayed in Fig. 1 for the $C$-parameter distribution. In this range the mean, median and mode for this sample are $0.1182$ for the first two and $0.1181$ for the mode. To a good approximation, we can thus describe the results obtained with a normal distribution of mean $\overline{\alpha}_{s}(M_{Z})=0.1182$ and variance $\sigma=0.0005$. We point out that while the minimum-$P_{\chi^{2}}$ strategy is applied to the entire range of values $10<C<$ 1, in order to isolate the normal behavior of the distribution, we must restrict the range to the perturbative region, where the theoretical predictions can be approximated by a linear model and the dominant source of error is statistics.
@@
Figure 1: A histogram of the result sample for the fitted values of $\alpha_{s}(M_{Z})$ for each possible interval in the range $0.720.38<C<$ 0.72 of the $C$-parameter, using PMC scale setting at $\sqrt{s}=M_{Z}$ center-of-mass energy. The normal distribution with mean $\overline{\alpha}_{s}(M_{Z})=0.1182$ and variance $\sigma=0.0005$ is also shown. The curve is normalized to the area of the histogram.
@@ IV Results
Results for the strong coupling extracted from thrust are shown in Tables 2, 3, 4 and from the $C$-parameter in Tables 5, 6, 7. We have labelled the different ranges (BFR), (TFR) and (GFR) to indicate respectively the Best Fitting Range, the Total Range of bins and the Range of bins used in Ref. [4]. In more detail, in the tables, we indicate with TFR the range of values $0.420.02<(1{-}T)<$ 0.42 for the thrust, excluding the first two bins, where the perturbative predictions need an improved description including resummation of the IR large logarithms and possibly also further nonperturbative or hadronization effects [7]. We underline that these corrections should be included also in the regions of the peak and in the multi-jet region above $(1{-}T)\simeq\frac{1}{3}$. We point out that in the multi-jet region we have also the kinematic constraint at LO that cancels the LO contribution and we thus expect that the missing higher-order contributions would lead to an improvement [5] in this region. For the same reasons we restrict the $50$-bin sample of the $C$-parameter to $46$ neglecting the first two bins below $C<0.04$ and the two bins at the border $C\sim 0.75$, where there is an enhancement of logarithm behavior and theoretical predictions need to be improved with the inclusion of resummation and of nonperturbative effects [6]. We label this range of values with (TFR1*). We use the label (TFR2*) to indicate the sample of bins where three bins below $C<0.06$ have been excluded, leading to a sample of $45$ bins.
@@
Analogous results are obtained in the other bin ranges. We note that the results shown in Table 2, for thrust using the PMC and the 5-loop solution for the running coupling have very small differences with respect to the particular range and that the same prediction is obtained for the coupling using the 5-loop perturbative solution of Eq. (I), as shown in Table 3 in the BFR range. It is to be noted that in the case of thrust results for extracted values of the coupling using the PMC in the BFR and in the total range of bins (TFR) are consistent within the errors. While for the $C$-parameter we find a significant reduction of the errors, of one order of magnitude in the range TFR1 using the PMC, independently from the particular solution used for the coupling as shown in Table 5 and Table 3. This effect is possibly due to an effect of constraint of the particular bin at $C\sim 0.05$. In fact, in the range of bins TFR2 that excludes this particular bin, the effect disappear. Given the poor statistics, we would tend to exclude this prediction, but in a global fit that includes also nonperturbative effects over the entire range of values (TFR1) of the $C$-parameter and that would improve the goodness of the fit in this range, we would expect a higher precision in the determination of the strong coupling. We show results using the conventional scale setting for thrust and the $C$-parameter in Table 4 and Table 7 respectively. Comparing the results to those shown in Ref. [4], we note that, by using the conventional scale setting, we obtain the same errors for the strong coupling extracted from the NNLO theoretical predictions of thrust and the $C$-parameter but with improved statistics. The extracted values of the strong coupling are analogous to those of Ref. [4]; although in the case of the $C$-parameter we find a small deviation. This is possibly due to the small differences in the NNLO coefficient that have also been shown in Ref. [37]. Comparison between conventional scale setting and PMC results clearly shows that PMC leads to a significantly higher precision.
@@
A further analysis has been performed to investigate possible biassing effect due to some correlations in the BFR strategy. For this purpose we have kept held the BFR range of bins set using the ALEPH data and then fitted the strong coupling at the $Z^{0}$ mass, $\alpha_{s}(M_{Z})$, using other independent data from the OPAL, DELPHI and L3 experiments [24, 25, 23]. Results are shown in Table 1.
@@
No particular bias effect due to the BFR approach is noted either for the central values or for the errors. The different central values obtained depend only on the experimental data used. Moreover, we find a reduced spread among the event-shape results according to the same experiment. The results appear to be in good agreement. Combining and performing the weighted average of these independent values, using the asymmetric quadratic model for the errors shown in Ref. [56] and including a correlation coefficient of $70\%$ between the event-shape variables thrust and $C$-parameter (as shown in Refs. [4, 22]), we obtain:
$\alpha_{s}(M_{Z})=0.1181^{+0.0005}_{-0.0005}.$ (10)
@@
Figure 3: Results for the strong coupling $\alpha_{s}(\mu)$ at the $Z^{0}$ peak for thrust and $C$-parameter in the BFR interval using ALEPH, OPAL, DELPHI and L3 [22, 24, 23, 25]. The shaded area shows the total errors for the average value $\alpha_{s}(M_{Z})=0.1181^{+0.0005}_{-0.0005}$.
In Fig. 3 we show the strong coupling $\alpha_{s}(M_{Z})$ result determined by the weighted average and also the results for $\alpha_{s}(M_{Z})$ from thrust and the $C$-parameter. Results for both thrust and the $C$-parameter are consistent with the world average [1] and also with recent PMC applications presented in Refs. [57, 58, 59, 60]. We point out that in this analysis we have not included part of the systematic errors, in particular those related to the hadronization effects. Thus, a direct comparison of the total error in Eq.10, with the total error obtained by summing in quadrature the statistical and theoretical errors (not the systematic error) obtained in Ref. [4], (i.e. $\Delta\alpha_{s}(M_{Z})\simeq\pm 0.0036$), shows that the PMC method can lead to a clear improvement of the precision with respect to the conventional scale setting adopted in Ref. [4].

### v4

@@ Abstract
We present a new approach to determining the strong coupling $\alpha_{s}(Q)$, over the entire range of validity of perturbative QCD, for scales above $\Lambda_{\mathrm{QCD}}$ and up to the Planck scale $\sim 1.22\cdot 10^{19}$ GeV, with the highest precision and using the data of a single experiment. In particular, we use the results obtained for the thrust ($T$) and $C$-parameter ($C$) distributions in $e^{+}e^{-}$ annihilation at a single annihilation energy $\sqrt{s}=M_{Z}$ (i.e. at the $Z^{0}$ peak). This new method is based on the intrinsic conformality (iCF) and on the Infinite-Order Scale Setting, using the Principle of Maximum Conformality (i.e. the PMC), which allows a rigorous determination of the renormalization scales for the event-shape variable distributions satisfying all of the requirements of Renormalization Group Invariance, including renormalization-scheme independence and consistency with Abelian theory in the $N_{C}\to 0$ limit. This new method is based on the scale-invariance of the iCF, which allows determination of $\alpha_{s}(\mu_{0})$ at any scale $\mu_{0}$, and on the Maximum Likelihood statistical approach.
We propose a novel approach to determining the best-fitting range by considering all possible intervals over the entire range of bins available in the perturbative region and selecting that which returns the most-likely-lowest $\chi^{2}_{\rm min}$. This new method is designed to eliminate the errors that arise due to selection of the bin-interval and that have been neglected in previous analyses. In particular, using data for thrust and $C$-parameter at the $Z^{0}$ peak from ALEPH, OPAL, DELPHI and L3 experiments, we obtain the average value: $\alpha_{s}(M_{Z})=0.1182^{+0.0007}_{-0.0007}$, for the strong coupling. This determination of $\alpha_{s}(M_{Z})$ is consistent with the world average and has an improved precision with respect to the values obtained from the analysis of event shape observables currently used in the world average.
@@
In this article we shall demonstrate that a recently proposed method [3] for solving the renormalization scale and scheme ambiguities, namely infinite-order scale setting using the principle of maximum conformality (PMC), based on intrinsic conformality (iCF) and by using the PMC, leads to an improved extraction of $\alpha_{s}(Q)$ at the $Z^{0}$ peak from the thrust and $C$-parameter distributions in the annihilation process $e^{+}e^{-}\to 3$ jets. Moreover, our new results agree with the most recent world average given by the Particle Data Group [1], improving the precision and significantly reducing the theoretical errors related to the renormalization scale.
In addition, we propose a new method to eliminate errors related to the choice of the bin-interval in the perturbative region, which is consistent with the Maximum-Likelihood approach and the $\chi^{2}$-minimization procedure used for the fit of the strong coupling.
@@ I The strong-coupling evolution
with the $\beta$-coefficients, $\beta_{n}$, n=0,,4n=0,...,4, from one- to five-loop accuracy respectively, calculated in a given scheme. Results for these coefficients can be found in Refs. [39, 40, 41, 42, 43, 44, 45, 46]. For a review of the strong coupling and of all the $\beta$ coefficients, see e.g. [47, 48]. In particular, we use the $\overline{\mathrm{MS}}$ scheme throughout this paper. For the running coupling $\alpha_{s}(Q)$, we use the solutions at 5-loop accuracy given by both the RunDec program [49] and by the perturbative solution given by
@@ II Event-shape variables and the renormalization scale
where $\vec{p}_{i}$ indicates the three-momentum of particle $i$ and the sums run over all particles in the hadronic final state; $\vec{n}$ is a unit vector that is varied to maximize thrust $T$, determining the so-called thrust axis, i.e. $\vec{n}_{T}$. Alternatively, the variable (1T)(1{-}T) is often used, which for 3-jet production at LO is restricted to the range $0<(1{-}T)<\frac{1}{3}$. We have back-to-back and spherically symmetric events for $T=1$ and $T=\frac{2}{3}$ respectively. For the $C$-parameter, $\theta_{ij}$ is the angle between $\vec{p_{i}}$ and $\vec{p_{j}}$. For 3-jet production at LO the $C$-parameter is restricted by kinematics to the range $0\leq C\leq\frac{3}{4}$.
Theoretical results for normalized infrared-safe observables, such as the thrust or $C$-parameter distribution for the annihilation process $e^{+}e^{-}\rightarrow 3\,$jets at the $Z^{0}$ peak to NNLO accuracy, are given in Refs. [50, 51, 33, 34, 35, 36, 37]. When using the conventional scale-setting method, only one value of $\alpha_{s}$, at the kinematic scale $\sqrt{s}$ of the process, can be extracted with errors given by the uncertainty in choosing the renormalization scale, while using the PMC, which is based on intrinsic conformality (iCF), it is possible to determine the strong coupling at any scale, even up to the highest scale where the theory is expected to breakdown. As shown in Eq. (I), the final renormalization scale $\mu$ is set by using the PMC method while $\mu_{0}$ can be set to any value in the perturbative range; we can thus determine $\alpha_{s}(\mu_{0})$ at any scale. Results for the PMC for both thrust and $C$-parameter have been discussed in Refs. [2, 52, 3, 53]. We remark that the PMC solves the renormalization scale ambiguity and the scales, related to each order of calculation, are fixed perturbatively by determining the $\beta_{0}$ term of each scale at each order of accuracy. These scales take into account the dynamics of the underlying quark and gluon subprocess and thus fix the magnitude of the strong coupling at each order and for the corresponding value of the selected variable (C,1T))(C,1{-}T)). In fact, these scales are not single valued but are themselves functions of the kinematic scale, $\sqrt{s}$, and of the event-shape variable via the coefficients of the number of flavors, $N_{f}$, related to the loop integrals and thus to their UV-divergent part. The PMC scales are fixed order by order and the last scale is set to the kinematic scale of the process $\sqrt{s}\equiv M_{Z}$, given the missing knowledge of the higher-order $\beta_{0}$-term. We have defined this type of uncertainty as the second kind of residual scale dependence [54]. We remark that both the first and second kind of residual scale uncertainties are of the same order as the uncertainties in the first missing higher-order contribution and thus both have a mere perturbative nature linked to the perturbative series and its accuracy.
@@ III Fitting Method
The $\chi^{2}$ analysis has been performed following the approach of Ref. [55]; thus, in the first approximation, considering the bins as being independent of each other and the experimental errors $\sigma_{i}$ determined by summing the statistical and systematic errors (due to detector effects and to hadronization) in quadrature. Effects due to the bin-to-bin correlation and to the intrinsic correlation among the event-shape results within one experiment, have been considered only in the final stage of the analysis, when all results are combined in the weighted average to obtain the highest precise determination of the strong coupling $\alpha_{s}(M_{Z})$.
@@
The high-precision experimental data have been taken from LEP [22]; in particular, these data consist of $42$ experimentally measured values for thrust, 0<(1T)<0.420<(1{-}T)<0.42, divided into $42$ bins centered and with spacing $0.01$, while for the $C$-parameter we have $50$ experimental measurements in the range $0<C<1$ divided in $50$ bins centered and with spacing of $0.02$. In order to fit the theoretical distributions to the experimental data, for the case of the $C$-parameter we have to reduce the $100$-bin partition shown in Ref. [37] to $50$ bins. This reduction is made by averaging over two consecutive values of the distributions for the LO, NLO and NNLO coefficients.
@@ III.1 Best-fitting region
The purpose of this analysis is to improve or at least to encode the method used in Ref. [4] of setting the fitting interval in the range where: “…  a good perturbative description is available and where hadronization and detector corrections are below …”, in order to eliminate any possible human prejudice or uncertainty due to the choice of the bin-interval. For the case of the (1T)(1{-}T) and $C$-parameter event-shape variables, we know well that in the central region of 3-jet production the perturbative calculations give the best fit to data. Besides, introduction of other corrections, such as nonperturbative corrections and hadronization over the entire range of the event shape, does not improve the theoretical predictions; on the contrary, much larger errors are introduced. Thus, it is a common practice to reduce the fitting region to the perturbative region in order to improve precision and statistics. Although we agree with this procedure, we must admit that it is not entirely consistent with the Maximum-Likelihood procedure adopted to fit the parameters. We are basically adding one degree of freedom – the interval range – and we are eliminating it subsequently by introducing some criteria, which might not be entirely consistent with the minimum $\chi^{2}$ analysis. Here, we now introduce a new procedure to select the BFR interval.
@@
Results of the fits for $\alpha_{s}(M_{Z})$, for each bin interval in the range $0.38<C<0.72$ (labelled as GFR) are displayed in Fig. 1 for the $C$-parameter distribution. In this range the mean, median and mode for this sample are $0.1182$ for the first two and $0.1181$ for the mode. To a good approximation, we can thus describe the results obtained with a normal distribution of mean $\overline{\alpha}_{s}(M_{Z})=0.1182$ and variance $\sigma=0.0005$. We point out that while the minimum-$P_{\chi^{2}}$ strategy is applied to the entire range of values $0<C<1$, in order to isolate the normal behavior of the distribution, we must restrict the range to the perturbative region, where the theoretical predictions can be approximated by a linear model and the dominant source of error is statistics.
@@
Figure 1: A histogram of the result sample for the fitted values of $\alpha_{s}(M_{Z})$ for each possible interval in the range $0.38<C<0.72$ of the $C$-parameter, using PMC scale setting at $\sqrt{s}=M_{Z}$ center-of-mass energy. The normal distribution with mean $\overline{\alpha}_{s}(M_{Z})=0.1182$ and variance $\sigma=0.0005$ is also shown. The curve is normalized to the area of the histogram.
@@ IV Results
Results for the strong coupling extracted from thrust are shown in Tables 3, 4, 5 and from the $C$-parameter in Tables 6, 7, 8. We have labelled the different ranges (BFR), (TFR) and (GFR) to indicate respectively the Best Fitting Range, the Total Range of bins and the Range of bins used in Ref. [4]. In more detail, in the tables, we indicate with TFR the range of values 0.02<(1T)<0.420.02<(1{-}T)<0.42 for the thrust, excluding the first two bins, where the perturbative predictions need an improved description including resummation of the IR large logarithms and possibly also further nonperturbative or hadronization effects [7]. We underline that these corrections should be included also in the regions of the peak and in the multi-jet region above $(1{-}T)\simeq\frac{1}{3}$. We point out that in the multi-jet region we have also the kinematic constraint at LO that cancels the LO contribution and we thus expect that the missing higher-order contributions would lead to an improvement [5] in this region. For the same reasons we restrict the $50$-bin sample of the $C$-parameter to $46$ neglecting the first two bins below $C<0.04$ and the two bins at the border $C\sim 0.75$, where there is an enhancement of logarithm behavior and theoretical predictions need to be improved with the inclusion of resummation and of nonperturbative effects [6]. We label this range of values with (TFR1*). We use the label (TFR2*) to indicate the sample of bins where three bins below $C<0.06$ have been excluded, leading to a sample of $45$ bins.
@@
Analogous results are obtained in the other bin ranges. We note that the results shown in Table 3, for thrust using the PMC and the 5-loop solution for the running coupling have very small differences with respect to the particular range and that the same prediction is obtained for the coupling using the 5-loop perturbative solution of Eq. (I), as shown in Table 4 in the BFR range. It is to be noted that in the case of thrust results for extracted values of the coupling using the PMC in the BFR and in the total range of bins (TFR) are consistent within the errors. While for the $C$-parameter we find a significant reduction of the errors, of one order of magnitude in the range TFR1 using the PMC, independently from the particular solution used for the coupling, as shown in Table 6 and Table 4. This effect is possibly due to an effect of constraint of the particular bin at $C\sim 0.05$. In fact, in the range of bins TFR2 that excludes this particular bin, the effect disappear. Given the poor statistics, we would tend to exclude this prediction, but in a global fit that includes also nonperturbative effects over the entire range of values (TFR1) of the $C$-parameter and that would improve the goodness of the fit in this range, we would expect a higher precision in the determination of the strong coupling. We show results using the conventional scale setting for thrust and the $C$-parameter in Table 5 and Table 8 respectively. Comparing the results to those shown in Ref. [4], we note that, by using the conventional scale setting, we obtain the same errors for the strong coupling extracted from the NNLO theoretical predictions of thrust and the $C$-parameter. The extracted values of the strong coupling are analogous to those of Ref. [4]; although in the case of the $C$-parameter we find a small deviation. This is possibly due to the small differences in the NNLO coefficient that have also been shown in Ref. [37]. Comparison between conventional scale setting and PMC results clearly shows that PMC leads to a significantly higher precision.
@@
A further analysis has been performed to investigate possible biassing effect due to some correlations in the BFR strategy. For this purpose we have kept held the BFR range of bins set using the ALEPH data and then fitted the strong coupling at the $Z^{0}$ mass, $\alpha_{s}(M_{Z})$, using other independent data from the OPAL, DELPHI and L3 experiments [24, 25, 23]. Results are shown in Table 1 and Table 2.
@@
No particular bias effect due to the BFR approach is noted either for the central values or for the errors. The different central values obtained depend only on the experimental data used. Moreover, we find a reduced spread among the event-shape results according to the same experiment. The results appear to be in good agreement.
In order to determine the strong coupling, we have first corrected the experimental errors obtained from the minimum-$\chi^{2}$-fit for each event-shape measurement, by introducing the effects of the bin-to-bin correlation. The corrected errors are indicated as: Expt. Err. in Tables: 1 and 2. In particular, a robust estimation of the covariance matrix and of the bin-to-bin correlation parameter has been performed by using the same strategy as indicated in Ref. [56] and in Ref. [57], that suggests to set the correlation parameter $\rho_{\rm eff}$ in order to obtain a consistent $\chi^{2}$: i.e. $\chi^{2}(\rho_{\rm eff})\equiv n_{\rm dof}$, when the $\chi^{2}_{\rm min}<n_{\rm dof}$. On the other hand, when $\chi^{2}_{\rm min}>n_{\rm dof}$, we rescale the errors until $\chi^{2}=n_{\rm dof}$ is satisfied. More in detail, we introduce the positive correlation for thrust and C-parameter results from ALEPH and for thrust results from the L3 experiment.111 We point out that this approach has some limitations, given by the assumption that the entire statistical gap is compensated with correlation.
Thus, combining and performing the weighted average of these independent values, using the asymmetric quadratic model for the errors shown in Ref. [58] and including also the correlation between the event-shape variables thrust and $C$-parameter using a correlation parameter of $70\%$ (as shown in Refs. [4, 22]), we obtain:
$\alpha_{s}(M_{Z})=0.1182^{+0.0007}_{-0.0007}.$ (10)
@@
Figure 3: Results for the strong coupling $\alpha_{s}(\mu)$ at the $Z^{0}$ peak for thrust and $C$-parameter in the BFR interval using ALEPH, OPAL, DELPHI and L3 [22, 24, 23, 25]. The shaded area shows the total errors for the average value $\alpha_{s}(M_{Z})=0.1182^{+0.0007}_{-0.0007}$.
In Fig. 3 we show the strong coupling $\alpha_{s}(M_{Z})$ result determined by the weighted average and also the results for $\alpha_{s}(M_{Z})$ from thrust and the $C$-parameter. Results for both thrust and the $C$-parameter are consistent with the world average [1] and also with recent PMC applications presented in Refs. [59, 60, 61, 62].
We remark that experimental errors have been obtained by first summing in quadrature statistical and systematic errors ( related to detector and hadronization effects ) taken directly from the experimental data [22, 24, 23, 25]) and subsequently correcting for bin-to-bin correlation or rescaling of the errors. We notice that in the highest-correlation case ( i.e. ALEPH), we obtain $\rho_{\rm eff}\simeq 0.93$ for thrust and $\simeq 0.83$ for C-parameter), the corrected experimental errors obtained by introducing bin-to-bin correlations, (i.e. $\sim\pm 0,0023$ for thrust and $\sim\pm 0,0020$ for the C-parameter) agree with those obtained by summing the independent experimental errors related to statistics, detector and hadronization, shown in Refs. [4, 22], (i.e. $\sim\pm 0.0021$ $\sim\pm 0.0022$ for thrust and $\sim\pm 0.0018$ , $\sim\pm 0.0017$ for the C-parameter respectively). Note that these results come from the different bin-intervals chosen in Refs. [4, 22]. This indicates that in the worst case (that of very high correlation) the BFR method can lead to errors that are no larger than the errors obtained by a separate treatment of the statistical and systematic errors, as shown in the previous analyses. An improved treatment of the errors (especially those coming from the hadronization models, e.g. including the full correlation matrices as obtained from the Monte Carlo simulations) would be necessary in order to test the BFR method and verify its results to high accuracy. Thus, if we neglect the systematic errors for the time being, a direct comparison of the total error in Eq.10, with the total error obtained by summing in quadrature the statistical and theoretical errors obtained in Ref. [4], (i.e. $\Delta\alpha_{s}(M_{Z})\simeq\pm 0.0036$), shows that the PMC method can lead to a clear improvement in precision with respect to the conventional scale setting adopted in Ref. [4].